OpenStudy (anonymous):

Find the zeros of the equation... 4x^2-x=5+y?

OpenStudy (dumbcow):

y = 4x^2-x-5 = 0 Factor by grouping -5*4 = -20 -5+4 = -1 4x^2 +4x -5x -5 = 0 4x(x+1) -5(x+1) = 0 (x+1)(4x-5) = 0 x+1 = 0 --> x = -1 4x-5 =0 --> x = 5/4